Sol Summary – April 18th

Crew 297 Sol Summary Report 18-04-2024

Sol: 04

Summary Title: stanleY yelnatS tuanortsA

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

Today we opened our eyes to a welcomed Martian morning partially filled with clouds. These clouds proved to be fortuitous during the later EVAs in reducing sun glare on screens and the need for respiration during digging. Today’s EVA was a first for our robotic crewmate Case. Case, roughly the size of its distant Martian cousin, Sojourner, has the intention of helping us take useful Martian imagery for planning purposes of future crews. Today was also the first time we had four crewmates involved in the same EVA and the first true attempt at digging the NPS. Case was able to capture useful imagery of the surroundings, yet, it was not able to take off its metaphorical leash, needing to be monitored constantly and manually driven for a large portion of its adventure. The other two EVA crew mates had great success – digging a hole suitable for the NPS at a rate of 30 cubic feet per hour. Two hours had passed before the rover and related GSE were collected, the hole was filled up, and we started heading back home. When the crew finally recovered from the eventful EVAs they learned about the functionality of WIDGIT and how to incorporate it into their collaborative report writing workflows. FInally, a savory Japanese curry dish was prepared and served by Chef Sean to the delight of the crews’ taste buds.

Look Ahead Plan:

Tomorrow we will have a chock-filled day of EVAs from 0900 to 1600. The first EVA will look to locate the NPS that has thankfully been placed by Mission Support for us. After they successfully find the NPS, they will provide the favor to the second EVA crew of hiding the NPS. The second EVA crew will then attempt to find the NPS with enough time to spare for additional gamma-ray spectral readings of the Martian rocks in the surrounding area.

Anomalies in work:

Suit 5 (ventilation electrical connection): Investigated and likely fixed. To be further examined/tested in the coming days.
GreenHab automated fan: The GreenHab fan powers on when the GH temp reaches 94F, but then shortly thereafter turns off. It power cycles continuously but never reaches its expected rpm.
Suit 4 (collar cable ring missing): Fixed today, will be functionally tested during EVA tomorrow.
ScienceDome: We have noticed a foul chemical smell. Dispositioning.
Morning (data from 0900): calm winds (6 mph), chilly (57F), and partially cloudy
Noon: calm winds (3 mph), warm (67F), and sunny
Afternoon (1800): calm winds (8 mph), warm (74F), and broken cloud coverage
Crew Physical Status: Crew health is nominal.

EVA: Nominal EVA. The crew of four involved with the EVA was able to test digging capabilities within their spacesuits and use the rover, Case, to capture MDRS site imagery.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (1 EVA)
EVA request (1 EVA)
Support Requested: We have requested that the [inert] NPS be placed by Mission Support at a location of their choosing (though reachable by rover and within a maximum of 15 minutes of walking) in either Candor Chasma, Eos Chasma, Valles Marineris, or Aurorae Chaos for a search and recovery EVA #6 on 19-04-2024. The NPS will be placed in the rear airlock tomorrow night, 18-04-2024 by 1600. Once the inert NPS is hidden, precise GPS coordinates and a photo are requested to be taken to not lose track of the location. These GPS coordinates are then asked to be sent to Pawel.

We also request to have the permission for Opportunity to be plugged in after EVA #6 tomorrow for one hour (1200-1300) between EVAs to extend its range.

Copyright © The Mars Society. All rights reserved. | Main Site